# Revisiting Robust RAG: Do We Still Need Complex Robust Training in the Era of Powerful LLMs?

### v1

@@ 4 Does Sophisticated Robust Training Still Matter in Powerful Models?
4.1 Do Sophisticated Document Selection Strategies Matter?
We conduct experiments to analyze the effectiveness of complex document selection strategies under Llama model families (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) in Table 1, and Qwen model families in Appendix Table 4.
@@ 4.1 Do Sophisticated Document Selection Strategies Matter?
The experimental results presented in Tables 1 and 4 provide compelling evidence that robust training significantly improves model resilience when processing noisy documents. While base models exhibit substantial performance degradation when encountering noisy documents during inference, models that undergo robust training maintain consistent and superior QA performance across various document selection strategies. A notable example is the Llama-2-7b-chat-hf model’s performance (Table 1) on the HotpotQA dataset, where training with golden documents improves the EM score from 3.3 (Base Model) to 30.67 (Golden Doc), indicating increased resilience to document noise. This pattern of improvement is consistently observed across both Llama (Table 1) and Qwen (Table 4) model families, strongly indicating that robust training effectively mitigates the base models’ inherent vulnerability to noisy documents.
@@
Experimental results indicate that the performance gains from sophisticated document selection strategies diminish as models evolve. For instance, comparing Llama-2-7b-chat-hf with Llama-3-8B-Instruct, the improvement in performance due to advanced document selection strategies decreases significantly, with the $\Delta$ (Worst $\rightarrow$ Best) metric for NQ dropping from 21.87% to 7.90% EM. A similar trend is observed when comparing Qwen1.5-7B-Chat to Qwen2.5-7B-Instruct, where the performance improvement from sophisticated document selection also shows a noticeable reduction. These results suggest that as models become more advanced, their ability to process and utilize information improves independently of complex document selection strategies, leading to diminished returns from such methods.
4.2 Do Adversarial Loss Functions Matter?
To investigate whether the design of complex adversarial loss functions contributes to model performance, in Table 1 and 4, we also analyze the robustness of various adversarial loss designs.
@@ 4.2 Do Adversarial Loss Functions Matter?
Figure 2: Comparison of F1 Scores on the HotpotQA dataset for training with Random Doc and Golden Doc across models with varying parameter sizes from 0.5B to 70B. We provide detailed results in Appendix Table 6.
@@
For the stronger models, such as Llama-3-8B-Instruct and Qwen2.5-7B-Instruct, the benefits of adversarial loss functions are less pronounced. The basic Llama-3-8B-Instruct already achieves an average EM / F1 of 34.78 / 45.97, and the introduction of adversarial losses (RAAT and IRM) results in only modest improvements, with average EM / F1 scores of 44.31 / 54.44 and 46.86 / 58.00, respectively. Similarly, Qwen2.5-7B-Instruct shows similar trends, with marginal gains from adversarial loss functions. These improvements are comparable to or even slightly worse than the performance achieved by random document selection (48.30 / 58.92) or top-1 document strategies (49.15 / 59.02). This indicates that the models’ inherent robustness reduces the impact of adversarial losses, and in some cases, may even hinder performance. We hypothesize that when the model’s internal robustness is already well-developed, additional constraints from adversarial losses may interfere with its ability to optimize on clean and relevant inputs.
@@
4.3 Do Training Strategies Matter Across Model Scales?
@@
Through analyzing Tables 1 and 4, we observe a counter-intuitive phenomenon: models trained with random documents outperformed those with golden documents, despite the latter containing ground truth answers typically yielding optimal results in standard SFT. To investigate whether this phenomenon extends beyond 7B-8B models, we conduct experiments across model scales from 0.5B to 70B parameters.
@@
The results (detailed in Appendix Table 6) in Figure 2 demonstrate that for smaller models ($\leq$3B parameters), training with golden documents leads to superior performance. This suggests that smaller models, limited by their inherent capabilities, benefit more from high-quality golden documents containing direct answers. However, as model size increases, we observe that training with random documents becomes more effective. This shift can be attributed to larger models’ enhanced question-answering abilities and improved robustness. These models can better generalize to downstream tasks even when trained on random documents, which may contain noisier or less structured information. This finding indicates that sophisticated document selection strategies become less crucial as model size increases, revealing an important scaling property in model training.
5 Why Sophisticated Training No Longer Matters in Powerful Models?
@@
Figure 3: Confidence scores for correct and wrong answers on HotpotQA dataset, comparing Llama2 and Llama3 models across various robust training methods.
5.1 Powerful Models Enable Natural Calibration
To understand whether powerful models inherently possess the ability to distinguish reliable from unreliable answers, we take HotpotQA dataset as an example to examine the confidence calibration capabilities for Llama2 (Llama-2-7b-chat-hf) and Llama3 (Llama-3-8B-Instruct) models. Here, confidence is the mean of token-wise probabilities in the model’s generated answer, providing a measure of the model’s certainty in its predictions Varshney et al. (2023); Xiong et al. (2024). Figure 3 reveals striking differences in their calibration patterns. In the base model, Llama2 shows poor natural calibration levels, where confidence scores for incorrect answers (95.8) abnormally exceed those for correct ones (93.8). In contrast, Llama3 demonstrates inherently better calibration, maintaining higher confidence for correct answers (97.5) than incorrect ones (91.3) without any specialized complex training.
While robust training methods (Golden Doc, Random Doc, and IRM) can effectively calibrate confidence scores and improve the gap between correct and incorrect answers for Llama2 from -2 to 12 with IRM, the marginal benefits of these complex training strategies diminish as model architectures advance. For Llama3, which already achieves a 6.2 confidence gap naturally, the improvements from these training methods become less significant. This finding strongly suggests that advances in model architecture can effectively eliminate the need for complex robustness training procedures, as newer models come with better built-in calibration capabilities.
Figure 4: Generalization performance comparison across different strategies trained on HotpotQA (diagonal hatches bars) and evaluated on NQ, WebQuestions, TriviaQA datasets (plain bars). More results available in Appendix Table 8.
5.2 Simple Training Strategies Generalize Well in Powerful Models
@@
As shown in Figure 4, simple strategies demonstrate surprisingly strong generalization ability. Random document selection matches or even outperforms sophisticated IRM across all evaluation datasets, with performance gaps of less than 1%. For instance, in TriviaQA, random selection (69.5 F1) slightly surpasses both golden (68.2 F1) and IRM (68.7 F1) approaches. This trend becomes more pronounced in the powerful Llama-3-8B-Instruct, where the performance gap between simple and sophisticated strategies further narrows. The consistent cross-dataset performance, regardless of training strategy, indicates that model capacity, rather than training sophistication, is the key driver of generalization ability. These findings provide strong evidence that as models become more powerful, sophisticated training strategies become increasingly unnecessary.
Figure 5: Attention visualization for a QA case. Each subplot shows attention distribution heatmaps across different models, where cell ($i$, $j$) represents the average attention weight from the $j$-th attention layer to document $i$. Text highlighted in green indicates the correct answer and corresponding Doc1, blue indicates key terms from the query, and red indicates incorrect model predictions. The color intensity in the heatmaps indicates attention strength.
5.3 Powerful Models Learn Effective Attention Patterns with Simple Training
@@
Figure 6: Performance comparison with different numbers of random documents during training. Increasing the number of random documents consistently improves model performance.
Figure 7: Training curves comparison between random and golden document strategies using Llama-2-7b-chat-hf and LLama-3-8B-Instruct.
5.4 Training with Random Docs: Better Performance and Faster Convergence

### v2

@@ 4. Diminishing Benefits of Complex Methods With Model Capacity
4.1. Marginal Benefits of Sophisticated Document Selection
We conduct experiments to analyze the effectiveness of complex document selection strategies under Llama model families (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) in Table 2 and Qwen model families (Qwen1.5-7B-Chat and Qwen2.5-7B-Instruct) in Table 3.
@@ 4.1. Marginal Benefits of Sophisticated Document Selection
The experimental results provide compelling evidence that robust training significantly improves model resilience when processing noisy documents. While base models exhibit substantial performance degradation when encountering noisy documents during inference, models that undergo robust training maintain consistent and superior QA performance across various document selection strategies. A notable example is the Llama-2-7b-chat-hf model’s performance (Table 2) on the HotpotQA dataset, where training with golden documents improves the EM score from 3.3 (Base Model) to 30.67 (Golden Doc), indicating increased resilience to document noise. This pattern of improvement is consistently observed across both Llama (Table 2) and Qwen (Table 3) model families, strongly indicating that robust training effectively mitigates the base models’ inherent vulnerability to noisy documents.
@@
Experimental results indicate that the marginal robust benefit from sophisticated document selection strategies diminish as models evolve. For instance, comparing Llama-2 with Llama-3, the improvement in performance due to advanced document selection strategies decreases significantly, with the $\Delta$ (Worst $\rightarrow$ Best) metric for WebQuestions dropping from 59.60% to 16.94% EM. A similar trend is observed when comparing Qwen1.5 to Qwen2.5, where the performance improvement from sophisticated document selection also shows a noticeable reduction. These results suggest that as models become more advanced, their ability to process and utilize information improves independently of complex document selection strategies, leading to diminished returns from such methods.
4.2. Marginal Benefits of Complex Loss Design
To investigate whether the design of complex adversarial loss functions contributes to model performance, in Table 2 and 3, we also analyze the robustness of various adversarial loss designs.
@@ 4.2. Marginal Benefits of Complex Loss Design
Figure 2. Comparison of F1 Scores on the HotpotQA (left) and NQ (right) dataset for training with Random Doc and Golden Doc across models with varying parameter sizes from 0.5B to 70B.
@@
For the stronger models, such as Llama-3-8B-Instruct and Qwen2.5-7B-Instruct, the benefits of adversarial loss functions are less pronounced. The basic Llama-3-8B-Instruct already achieves an average EM / F1 of 34.78 / 45.97, and the introduction of adversarial losses (RAAT and IRM) results in only modest improvements, with average EM / F1 scores of 44.46 / 54.55 and 46.97 / 58.16, respectively. This indicates that the models’ inherent robustness reduces the impact of adversarial losses, and in some cases, may even hinder performance. We hypothesize that when the model’s internal robustness is already well-developed, additional constraints from adversarial losses may interfere with its ability to optimize on clean and relevant inputs.
Conclusion: the necessity of adversarial loss.
@@
4.3. Scaling Behavior and Cross-Domain Generalization
In this section, we conduct experiments across a wide range of model scales and diverse domains to investigate how the effectiveness of robust RAG training strategies varies with model capabilities.
4.3.1. Parameter Scale Analysis
@@
Through analyzing Tables 2 and 3, we observe a counter-intuitive phenomenon: models trained with random documents outperformed those with golden documents, despite the latter containing ground truth answers typically yielding optimal results in standard SFT. To investigate whether this phenomenon extends beyond 7B-8B models, we conduct experiments across model scales from 0.5B to 70B parameters.
@@
The results in Figure 2 demonstrate that for smaller models ($\leq$3B parameters), training with golden documents leads to superior performance. This suggests that smaller models, limited by their inherent capabilities, benefit more from high-quality golden documents containing direct answers. However, as model size increases, we observe that training with random documents becomes more effective. This shift can be attributed to larger models’ enhanced question-answering abilities and improved robustness. These models can better generalize to downstream tasks even when trained on random documents, which may contain noisier or less structured information. This finding indicates that sophisticated document selection strategies become less crucial as model size increases, revealing an important scaling property in model training.
4.3.2. Cross-Domain Generalization
To validate our findings beyond general question-answering tasks, we examine whether this phenomenon generalizes across temporal reasoning (TimeQA (Chen et al., 2021)), specialized legal knowledge (LegalBench (Guha et al., 2023)), and scenarios with conflicting evidence (RAGuard (Zeng et al., 2025)).
Model Method TimeQA
EM F1
Llama-2-7b-chat-hf Golden Doc 32.93 43.44
Top-1 Doc 32.26 43.13
Random Doc 31.90 43.32
RetRobust 31.97 42.58
$\Delta$ (Worst $\rightarrow$ Best) 3.23% 2.02%
Llama-3-8B-Instruct Golden Doc 35.44 46.90
Top-1 Doc 35.27 46.94
Random Doc 35.87 47.20
RetRobust 35.90 47.15
$\Delta$ (Worst $\rightarrow$ Best) 1.79% 0.63%
Table 4. Performance comparison of different LLMs (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) across various robust RAG methods on the TimeQA dataset.
Temporal Reasoning Domain
We evaluate the performance of various document selection strategies on the TimeQA dataset (Chen et al., 2021), which specifically tests models’ ability to handle time-sensitive information. As shown in Table 4, we compare two different language models: Llama-2-7b-chat-hf and Llama-3-8B-Instruct. With Llama-2-7b-chat-hf, we observe moderate variations across strategies, with a delta of 3.23% in EM and 2.02% in F1 between best and worst approaches. Golden Document selection generally yields the strongest performance. However, with Llama-3-8B-Instruct, this marginal robustness benefit narrows significantly to just 1.79% for EM and 0.63% for F1, with RetRobust and Random Document selection slightly outperforming Golden Document selection.
Model Method LegalBench
Llama-2-7b-chat-hf Base Model 2.55
RALM 70.96
RetRobust 66.67
Top-1 Doc 71.21
Golden Doc 70.96
Random Doc 70.96
Irrelevant Doc 72.73
IRM 69.19
$\Delta$ (Worst $\rightarrow$ Best) 9.09%
Llama-3-8B-Instruct Base Model 73.23
RALM 85.61
RetRobust 84.85
Top-1 Doc 85.86
Golden Doc 86.62
Random Doc 86.87
Irrelevant Doc 86.36
IRM 83.84
$\Delta$ (Worst $\rightarrow$ Best) 3.61%
Table 5. Performance comparison of the Llama-2-7b-chat-hf and Llama-3-8B-Instruct models across various robust RAG methods on the Legal-Bench Consumer Contracts QA dataset.
Specialized Legal Domain
We extend our investigation to specialized domains by evaluating various robust RAG methods on the LegalBench Consumer Contracts QA dataset (Guha et al., 2023). Table 5 presents a comprehensive comparison across both Llama-2-7b-chat-hf and Llama-3-8B-Instruct models. The results strongly reinforce our core findings. For Llama-2-7b-chat-hf, we observe substantial performance variations across different methods, with a 9.09% delta between the best (Irrelevant Doc at 72.73%) and worst (RetRobust at 66.67%) effective approaches. This significant variation suggests that older models remain highly sensitive to the specific training methodology employed. In contrast, the more advanced Llama-3-8B-Instruct model demonstrates higher baseline performance (73.23% without retrieval augmentation) and reduced sensitivity to the choice of method. The performance delta between the best (Random Doc at 86.87%) and worst effective method (IRM at 83.84%) narrows to just 3.61% - less than half the gap observed with Llama-2.
Model Method RAGuard
Llama-2-7b-chat-hf
Random Doc 56.04
Top-1 Doc 56.73
Golden Doc 54.31
Irrelevant Doc 56.00
IRM 51.81
RetRobust 46.71
$\Delta$ (Worst $\rightarrow$ Best) 21.45%
Llama-3-8B-Instruct
Random Doc 59.82
Top-1 Doc 60.08
Golden Doc 59.37
Irrelevant Doc 59.14
IRM 58.95
RetRobust 59.78
$\Delta$ (Worst $\rightarrow$ Best) 1.92%
Table 6. Performance comparison of different LLMs (Llama-2-7b-chat-hf and Llama-3-8B-Instruct) across various robust RAG methods on the RAGuard Benchmark.
Evaluation with Conflicting Evidence
To validate our findings in challenging scenarios, we evaluate robust RAG methods on the RAGuard benchmark (Zeng et al., 2025), which tests model robustness against contradictory retrievals. Table 6 shows that for Llama-2-7b-chat-hf, there is a substantial 21.45% performance gap between the best and worst training methods. In contrast, Llama-3-8B-Instruct demonstrates remarkably consistent performance, with the gap shrinking to merely 1.92% - a 90% reduction compared to the less capable model. These results on a benchmark with conflicting evidence further confirm our central finding that the benefits of complex robust training methods diminish significantly as model capacity increases.
Across all three domains, we consistently observe that performance variations between different document selection strategies decreases significantly with more advanced models. These experiments confirm that as language models increase in capability, the benefits of complex robust RAG training strategies diminish significantly, regardless of domain specificity.
5. Understanding Diminishing Benefits of Complex Training in Powerful LLMs
@@
Figure 3. Confidence scores for correct and wrong answers on HotpotQA dataset, comparing Llama2 and Llama3 models across various robust training methods.
5.1. Powerful Models Enable Natural Calibration
To understand whether powerful models inherently possess the ability to distinguish reliable from unreliable answers, we take HotpotQA dataset as an example to examine the confidence calibration capabilities for Llama2 (Llama-2-7b-chat-hf) and Llama3 (Llama-3-8B-Instruct) models. Here, confidence is the mean of token-wise probabilities in the model’s generated answer, providing a measure of the model’s certainty in its predictions (Varshney et al., 2023; Xiong et al., 2024). Figure 3 reveals striking differences in their calibration patterns. In the base model, Llama2 shows poor natural calibration levels, where confidence scores for incorrect answers (95.8) abnormally exceed those for correct ones (93.8). In contrast, Llama3 demonstrates inherently better calibration, maintaining higher confidence for correct answers (97.5) than incorrect ones (91.3) without any specialized complex training.
While robust training methods can effectively calibrate confidence scores and improve the gap between correct and incorrect answers for Llama2 from -2 to 12 with IRM, the marginal benefits of these complex training strategies diminish as model architectures advance. For Llama3, which already achieves a 6.2 confidence gap naturally, the improvements from these training methods become less significant. This finding strongly suggests that advances in model architecture can effectively eliminate the need for complex robustness training procedures, as newer models come with better built-in calibration capabilities.
Figure 4. Generalization performance comparison across different strategies trained on HotpotQA (diagonal hatches bars) and evaluated on NQ, WebQuestions, TriviaQA datasets (plain bars).
5.2. Simple Training Strategies Generalize Well in Powerful Models
@@
Figure 5. Attention visualization for a QA case. Each subplot shows attention distribution heatmaps across different models, where cell (ii, $j$) represents the average attention weight from the $j$-th attention layer to document ii. Text highlighted in green indicates the correct answer and corresponding Doc1, blue indicates key terms from the query, and red indicates incorrect model predictions. The color intensity in the heatmaps indicates attention strength.
As shown in Figure 4, simple strategies demonstrate surprisingly strong generalization ability. Random document selection matches or even outperforms sophisticated IRM across all evaluation datasets, with marginal robustness benefit of less than 1%. For instance, in TriviaQA, random selection (69.5 F1) slightly surpasses both golden (68.2 F1) and IRM (68.7 F1) approaches. This trend becomes more pronounced in the powerful Llama-3-8B-Instruct, where the performance gap between simple and sophisticated strategies further narrows. The consistent cross-dataset performance, regardless of training strategy, indicates that model capacity, rather than training sophistication, is the key driver of generalization ability. These findings provide strong evidence that as models become more powerful, sophisticated training strategies become increasingly unnecessary.
5.3. Powerful Models Learn Effective Attention Patterns with Simple Training
@@
Figure 6. Performance comparison with different numbers of random documents during training. Increasing the number of random documents consistently improves model performance.
Figure 7. Training curves comparison between random and golden document strategies using Llama-2-7b-chat-hf and LLama-3-8B-Instruct.
5.4. Training with Random Docs: Better Performance and Faster Convergence